# Survey of Graph Neural Network for Internet of Things and NextG Networks

### v1

@@ III GNN for Internet of Things
Internet of Things refers to a network of interconnected devices that can communicate and share data with each other over the internet, without requiring direct human interaction. These devices are embedded with sensors, software, and other technologies that enable them to collect and exchange data, as well as perform specific functions. There are two major challenges with IoT. First, due to the vast number of devices connected through the internet, it is important to safeguard the devices from adversaries that may infiltrate the network and second, for applications that use different types of sensors to keep track of different entities, it is important to collect and process the data more efficiently. Towards this goal, IoT Network Intrusion and IoT Data Fusion have been extensively studied. We discuss the applications of GNN for IoT Network Intrusion Detection in Section III-A and IoT Data Fusion in Section III-B. Figure 6 shows representative IoT sensors and intrusion attacks.
III-A IoT Network Intrusion Detection
IoT Network Intrusion Detection (NID) refers to the process of monitoring and analyzing the traffic and behavior within an IoT network to identify and respond to unauthorized or malicious activities. IoT devices are increasingly being integrated into various industries and applications, ranging from smart homes and industrial automation to healthcare and transportation. However, the proliferation of IoT devices also presents new security challenges and vulnerabilities, making it crucial to have effective intrusion detection mechanisms in place for safeguarding IoT ecosystems, as these devices often have limited computing resources and may lack the robust security features found in traditional computing devices. By implementing effective intrusion detection systems, organizations can better protect their IoT networks from various threats, including malware, unauthorized access, data breaches, and other security risks. Below we discuss a few interesting approaches proposed for IoT NID. These are divided into general intrusion detection applications [132, 177, 178, 179] and Malicious logins and Malware detection [180, 181].
In [132], the authors present a GNN-based approach called Edge-Graph SAmple and aggreGatE aka E-GraphSAGE that allows capturing both the edge features as well as the topological information for network intrusion detection in IoT networks. Compared to traditional GraphSAGE [131] which mainly focuses on node features for node classification but falls short on edge classification using edge features. Classification based on edge features allows us to classify different network flows to the nodes into benign and attack flows which can be achieved using E-GraphSAGE. Similarly, in [177], the authors propose Anomal-E, a GNN-based approach to detect both anomaly and intrusion in computer networks using edge features and a graph topological structure in a self-supervised manner. In [178], the authors present a general framework for intrusion detection called Graph Intrusion Detection designed based on GNN for industrial IoT applications that are highly susceptible to cyber-attacks. This framework is designed to learn both the network topology as well as the GNN network parameter at the same time. While there are several ML-based intrusion detection techniques, they are not robust to common adversarial attacks. To address this challenge, in [179] the authors present a novel GNN-based model designed to process and learn from graph-structured information that keeps flow records and their relationships compared to other techniques that only focuses only on the different flows between the source and the destination hosts.
Another possible network intrusion scenario is malicious logins and the use of malware to gain access to the network. In [180], to address the challenge of malicious login the authors present a GNN-based solution called MLTracer. This framework consists of two key features. First, it enables differentiating crucial attributes of logins such as the source/destination IDs, roles, port, and authentication type, among others. Second, a GNN model in conjunction with CNN is used to detect malicious logins. In [181], the authors present a Network Flow GNN for malware detection and classification. The framework extracts the flow graphs to classify them using a novel edge feature-based graph neural network model. Three derived model variants are presented namely a graph classifier, a graph autoencoder, and a one-class graph neural network.
@@ III-B IoT Data Fusion
IoT data fusion, also known as information fusion, sensor data fusion, or multi-sensor data fusion, refers to the process of integrating and combining data from multiple IoT devices and sensors to obtain a more comprehensive and accurate understanding of a physical or digital environment. The primary goal of IoT data fusion is to extract valuable insights, make informed decisions, and improve the overall efficiency and effectiveness of IoT systems. To implement IoT data fusion, various techniques and technologies can be used, including data preprocessing, sensor calibration, data alignment, feature extraction, ML algorithms [182, 183], and data visualization tools. The choice of fusion method depends on the specific use case and the characteristics of the data sources involved.
@@
Figure 8: Applications of GNN for RF spectrum sensing tasks.
IV GNN for Spectrum Awareness
As the number of wireless devices increases the spectrum becomes crowded where the same frequency band is shared among different services and applications. This causes interference and degradation in QoS which makes it critical to be aware of the spectrum to avoid interfering with licensed/unlicensed incumbent users and to enable effective use of the spectrum. To this end, many ML techniques have been used to enable spectrum awareness and GNN is one of the key enabling techniques due to its ability to model complex and dynamic environments. In this section, we shed light on the use of GNN for Radio Frequency Signal Classification (Section IV-A) and RF Spectrum Sensing (Section IV-B).
IV-A Radio Frequency Signal Classification
ML is known for its classification and prediction capabilities across domains. Traditionally, CNNs [184, 185, 186, 187, 188] and RNNs have dominated RF signal classification tasks such as modulation classification, signal type classification, RF Fingerprinting among others [188, 189, 190]. More recently, GNNs are also being leveraged for these tasks including but not limited to [191, 192]. Figure 7 shows common applications of GNN for RF signal classification and prediction tasks. Furthermore, GNN has also impacted other classification tasks such as anomaly classification [193] and traffic classification [194].
AvgNet [191] is an Adaptive Visibility Graph algorithm that maps time series data into graphs and can be used to classify various signal modulations. The main advantage of this algorithm is that it can automatically adapt into best-matching graphs, reducing the time consumed during the mapping process and aggregating both the local and global information of the signal sample. The proposed model is verified using simulation and the results show that AvgNet can effectively classify 11 modulations namely Amplitude Modulation - Double Sideband (AM-DSB), Amplitude Modulation - Single Sideband (AM-SSB), Phase Shift Keying (8PSK), Binary Phase Shift Keying (BPSK), Continuous Phase Frequency Shift Keying (CPFSK), Gaussian Frequency Shift Keying (GFSK), Pulse Amplitude Modulation (PAM4), Quadrature Amplitude Modulation (QAM16, QAM64), Quadrature Phase Shift Keying (QPSK) and Wideband Frequency Modulation (WBFM).
@@ IV-A Radio Frequency Signal Classification
GNN also finds its application for traffic classification scenarios where it is important to differentiate between traffic generated by various applications while identifying any encrypted or malicious traffic. To enable this, in [194], the authors present CGNN, a Chained GNN to effectively classify various real-world traffic scenarios and categorize them into the different applications as well as label them as ”normal”, ”encrypted” or ”malicious” traffic. Through extensive simulation, it is shown that compared to CNN and other state-of-the-art techniques, GNN, specifically CGNN outperforms them in terms of three key metrics namely precision, accuracy, and recall.
IV-B RF Spectrum Sensing

### v2

@@ III GNN for Internet of Things
Internet of Things refers to a network of interconnected devices that can communicate and share data with each other over the internet, without requiring direct human interaction. These devices are embedded with sensors, software, and other technologies that enable them to collect and exchange data, as well as perform specific functions. In Section III-A we first discuss the application of GNN for different IoT smart applications. There are two major challenges with IoT systems. First, due to the vast number of devices that are heterogeneous in nature, it is important to collect all the data and combine it to extract the relevant information. Second, it is also important to safeguard the devices from adversaries that may infiltrate the network and spoof or steal valuable data. Therefore it is important to address these two challenges. With this challenge as the motivation, IoT Data Fusion and IoT Network Intrusion have been extensively studied. Unlike CNNs, which work best on grid-like structures (e.g., images), and RNNs, which excel at sequential data (e.g., time series data), GNNs can process non-grid-like and non-sequential unstructured data where relationships are more complex and are not fixed. This makes them ideal for aggregating and using data from heterogeneous IoT applications. To this end, we discuss the applications of GNN for IoT Data Fusion in Section III-B and Network Intrusion Detection in Section III-C. Figure 9 shows representative IoT sensors and intrusion attacks. Table V provides a high-level comparison of surveyed paper using GNNs for IoT applications. We also provide the dataset or the simulator used for performance evaluation.
TABLE V: Comparison of surveyed papers using GNNs for Internet of Things applications.
Section Paper Application ML Algorithm Used Dataset Used
Section III-A IoT Smart Applications Li et al. [180] Human Pose Forecasting System Graph Convolutional Neural Network Human 3.6M dataset [181], Penn Action dataset [182]
Yanmei et al. [183] Short-term load prediction Graph Attention Network + LSTM Short-term power load data - Shijiazhuang power supply company
Zhang et al. [184] Crowd flow prediction Spatio-temportal GNN University of Macau Campus WiFi Data
Sharma et al. [185] Real-time traffic-speed estimation Spatio-Temportal Graph Attention Network PeMS Data [186]
Pamuklu et al. [187] Smart agriculture GNN-RL Simulated dataset
Section III-B IoT Data Fusion Ma et al. [188] Air quality prediction Spatio-Temporal Graph Convolutional Network Air Pollution and Meteorological Data [189]
Yang et al. [190] Pedestrian trajectory prediction Spatio-Temporal Graph Convolutional Network ETH [191], UCY [191]
Section III-C IoT Network Intrusion Detection Lo et al. [144] IoT intrusion detection GraphSAGE ToN-IoT [192]; BoT-IoT [193]; NF-TON-IoT and NF-BoT-IoT [194]
Caville et al. [195] Anomaly and intrusion detection GraphSAGE NF-UNSW-NB15-v2 [196]; NF-CSE-CIC-IDS2018-v2 [197]
Zhang et al. [198] Industrial IoT Intrusion Detection Graph Convolutional Network Industrial Control System Traffic Data Sets [199]
Pujol-Perich et al. [200] Network Intrusion Detection MPNN GNN CIC-IDS2017 [201]
Xiao et al. [122] Malicious insider threat detection Graph Capsule Networks Universal insider threat dataset CERT v4.2[202]
Liu et al. [203] Malicious login detection GNN+CNN Los Alamos National Lab’s (LANL’s) comprehensive cybersecurity events dataset [204], Own corporate dataset
Busch et al. [205] Malware detection and classification Graph Autoencoder CICAndMal2017 [206]
III-A IoT Smart Applications
In simple terms, IoT smart applications are those that leverage data gathered from the users to improve their experience and quality of life through the process of data analytics and use AI/ML for recommending insights based on user’s past activity. As shown in Figure 1 there are several IoT smart applications. In this section, we will focus on a subset of applications that leverage GNN for the data analytics and forecasting [180, 183, 184], real-time estimation and decision-making [185], efficient task offloading and planning [187]. In [180] the authors propose a GCN-based hybrid cloud-edge system called GPFS, a graph-based pose forecasting system that is capable of online learning that can significantly improve the forecasting accuracy in new environments. This system is specifically implemented for a smart home deployment. In [183] the authors propose a GAT-LSTM-based prediction algorithm called EnGAT-BiLSTM for load prediction in smart energy applications where GAT is used to extract high-quality load spatio-temporal features while bi-directional LSTM (BiLSTM) is used for lifelong learning of load prediction. CrowdTelescope [184] is a Wi-Fi positioning-based multi-grained spatio-temporal crowd flow prediction framework that leverages the extracted human mobility trace from the Wi-Fi connection records on the university campus which can then be used to predict the crowdflow within the campus with different granularities. In [185] the authors propose a Spatio-Temporal Graph Attention Network-based solution for real-time traffic-speed estimation for smart city applications by considering the structure and characteristics of road networks. In [187] the authors propose a GNN-RL-based solution to optimize the task offloading from IoT devices to the UAVs for smart agriculture applications. The objective of the solution is to reduce the task deadline violations and optimize the energy usage of the UAVs to extend their mission time.
@@ III-B IoT Data Fusion
In the previous section, we discussed several IoT smart applications that gather different types of data and how it is used for different applications. There is one important step within different IoT smart applications which is called IoT data fusion. IoT data fusion, also known as information fusion, sensor data fusion, or multi-sensor data fusion, refers to the process of integrating and combining data from multiple IoT devices and sensors to obtain a more comprehensive and accurate understanding of a physical or digital environment. The primary goal of IoT data fusion is to extract valuable insights, make informed decisions, and improve the overall efficiency and effectiveness of IoT systems. To implement IoT data fusion, various techniques and technologies can be used, including data preprocessing, sensor calibration, data alignment, feature extraction, ML algorithms [188, 190], and data visualization tools. The choice of fusion method depends on the specific use case and the characteristics of the data sources involved.
@@
III-C IoT Network Intrusion Detection
In the previous section we discussed the applications of GNN for IoT data fusion where data from different heterogeneous data are collected and analyzed for various applications such as surveillance, and autonomous driving among others. But the major challenge associated with these kinds of data generating and data storing applications is that they are susceptible to various attacks as shown in Fig. 9. These kinds of attacks are also called Network Intrusion attacks. It is important to design countermeasure approaches to overcome these challenges. These approaches are broadly called IoT Network Intrusion Detection approaches. IoT Network Intrusion Detection (NID) refers to the process of monitoring and analyzing the traffic and behavior within an IoT network to identify and respond to unauthorized or malicious activities. IoT devices are increasingly being integrated into various industries and applications, ranging from smart homes and industrial automation to healthcare and transportation. However, the proliferation of IoT devices also presents new security challenges and vulnerabilities, making it crucial to have effective intrusion detection mechanisms in place for safeguarding IoT ecosystems, as these devices often have limited computing resources and may lack the robust security features found in traditional computing devices. By implementing effective intrusion detection systems, organizations can better protect their IoT networks from various threats, including malware, unauthorized access, data breaches, and other security risks. Below we discuss a few interesting approaches proposed for IoT NID. These are divided into general intrusion detection applications [144, 195, 198, 200, 122] and Malicious logins and Malware detection [203, 205].
In [144], the authors present a GNN-based approach called Edge-Graph SAmple and aggreGatE aka E-GraphSAGE that allows capturing both the edge features as well as the topological information for network intrusion detection in IoT networks. Compared to traditional GraphSAGE [143] which mainly focuses on node features for node classification but falls short on edge classification using edge features. Classification based on edge features allows us to classify different network flows to the nodes into benign and attack flows which can be achieved using E-GraphSAGE. Similarly, in [195], the authors propose Anomal-E, a E-GraphSAGE-based approach to detect both anomaly and intrusion in computer networks using edge features and a graph topological structure in a self-supervised manner. In [198], the authors present a general framework for intrusion detection called Graph Intrusion Detection designed based on GNN for industrial IoT applications that are highly susceptible to cyber-attacks. This framework is designed to learn both the network topology as well as the GNN network parameter at the same time. While there are several ML-based intrusion detection techniques, they are not robust to common adversarial attacks. To address this challenge, in [200] the authors present a novel GNN-based model designed to process and learn from graph-structured information that keeps flow records and their relationships compared to other techniques that focuses only on the different flows between the source and the destination hosts. In [122], the authors propose CapsITD, a novel user-level threat detection method using GCapsNet. The GCapsNet is used to learn the statistical features of users’ daily activities and then use the learned features to detect any malicious activity and threats.
Another possible network intrusion scenario is malicious logins and the use of malware to gain access to the network. In [203], to address the challenge of malicious login, the authors present a GNN-based solution called MLTracer. This framework consists of two key features. First, it enables differentiating crucial attributes of logins such as the source/destination IDs, roles, port, and authentication type, among others. Second, a GNN model in conjunction with CNN is used to detect malicious logins. In [205], the authors present an autoencoder-based Network Flow GNN for malware detection and classification. The framework extracts the flow graphs to classify them using a novel edge feature-based graph neural network model. Three derived model variants are presented namely a graph classifier, a graph autoencoder, and a one-class graph neural network.
III-D Lessons Learned
In this section, we present the key takeaways and lessons learned from the applications of GNN for IoT applications discussed in Sections III-A - III-C.
  1. 1.
    Traditional ML models including deep learning architectures like CNNs fail to leverage the relational information present in IoT networks due to structured grid-like data whereas GNN are specifically designed to process graph-structured data, allowing them to capture complex inter-dependencies, dynamically update node representations based on their neighbors, and better model IoT communication patterns. This capability is crucial in IoT applications, where network effects, communication patterns, and distributed sensing play a significant role in data processing and decision-making.
  2. 2.
    As discussed earlier, GNNs propagate information across edges in a graph, aggregating and transforming node features. However, as the number of nodes and edges in an IoT network grows, the computational burden increases significantly. Each node’s representation is influenced by its neighbors, and in deep GNN architectures, this information propagates across multiple layers, leading to an exponential increase in computation. This scalability challenge makes it difficult to deploy GNNs in large-scale IoT networks, where thousands or even millions of devices may be interconnected.
  3. 3.
    Many IoT applications, such as autonomous driving, industrial automation, and smart healthcare, require real-time decision-making. GNNs, however, involve iterative message-passing mechanisms that increase inference time. Unlike simpler models that can quickly process incoming data, GNNs may struggle to meet the low-latency requirements of real-time IoT systems. This limitation necessitates optimizations such as model compression, approximate inference techniques, or hardware acceleration to make GNNs viable for real-time scenarios.
  4. 4.
    IoT ecosystems generate diverse data types such as temperature and humidity readings from environmental sensors, images from surveillance cameras, text-based logs from communication protocols, and more. GNNs must integrate and process this heterogeneous data efficiently while maintaining meaningful graph representations. Achieving this requires specialized architectures, such as multi-modal GNNs, that can incorporate different data modalities into a unified learning framework. Traditional ML models treat different data types separately and often require handcrafted feature fusion techniques, which are inefficient and inflexible. GNNs, however, naturally accommodate heterogeneous data by encoding different modalities into node and edge features. Models like Heterogeneous GNNs (H-GNNs) [207, 208] are designed specifically to handle multi-modal IoT data, making them far more adaptable and effective.
  5. 5.
    IoT systems are often deployed in open environments, making them susceptible to cyber threats such as data spoofing, node compromise, and adversarial attacks on machine learning models. GNNs, like other deep learning models, can be fooled by subtle perturbations in the input data, leading to incorrect predictions or system failures. Developing robust GNN architectures that can detect and mitigate adversarial attacks is essential for maintaining the security and reliability of IoT networks. Traditional ML models lack built-in robustness against adversarial threats because they assume independent and identically distributed (IID) data points. GNNs, however, can leverage their relational structure to detect anomalous patterns and use graph-based anomaly detection methods to identify malicious behavior. Additionally, techniques such as adversarial training [209, 210] and robust graph learning algorithms [211, 212] can enhance GNN’s resilience against security threats in IoT networks.
  6. 6.
    Many IoT devices run on battery power or energy-harvesting techniques, limiting their computational capacity. The high computational requirements of GNNs pose a challenge in such environments, as frequent model updates and inference operations can quickly deplete energy reserves. Techniques like model pruning, quantization, and edge computing strategies are needed to adapt GNNs for energy-efficient deployment in resource-constrained IoT settings. Additionally, deploying tiny GNNs [213, 214] on low-power IoT nodes can help reduce energy consumption while maintaining high performance.
  7. 7.
    While GNNs excel at capturing relational structures, they may not be ideal for all aspects of IoT data processing. For instance, CNNs are highly effective at processing image data, while RNNs and transformers are well-suited for sequential and time-series data. By integrating GNNs with these models, hybrid architectures can be developed to leverage the strengths of each, enhancing overall performance in IoT applications. Such integrations enhance the overall accuracy and efficiency of IoT analytics.
  8. 8.
    Despite the growing interest in applying GNNs to IoT networks, there is a lack of widely accepted benchmarks tailored to this domain. Standardized datasets and evaluation protocols are crucial for comparing different GNN architectures, measuring their real-world effectiveness, and fostering innovation. Traditional ML models have well-established datasets, but GNNs in IoT lack such standardized resources. Creating open-source IoT-specific datasets incorporating factors such as device mobility, communication patterns, and heterogeneous data streams would enable more rigorous testing and validation of GNN-based solutions.
  9. 9.
    In conclusion, GNNs offer powerful tools for IoT applications, particularly in dealing with complex, interconnected systems. The future of GNNs in IoT lies in optimizing these models for practical, real-world deployment, ensuring they are both effective and efficient in a variety of IoT scenarios.
IV GNN for Spectrum Awareness
As the number of wireless devices continues to surge, the radio spectrum is becoming increasingly congested, leading to unwanted interference and a decline in Quality of Service (QoS). This challenge is particularly critical as modern communication systems must coexist within shared frequency bands, often competing with both licensed and unlicensed users. Without effective spectrum awareness, networks risk inefficient spectrum utilization and disruptive interference, ultimately degrading performance for essential applications. To tackle this challenge, advanced ML techniques have been leveraged to enhance spectrum awareness, enabling intelligent and adaptive spectrum management. GNNs have emerged as a powerful tool due to their unique ability to model complex, dynamic spectral environments. Unlike traditional ML approaches, GNNs can effectively capture spatial and relational dependencies among spectrum users, making them well-suited for tasks such as spectrum sensing, and adaptive signal classification. To this end, in this section, we shed light on the use of GNN for Radio Frequency Spectrum Sensing in Section IV-A and Radio Frequency Signal Classification in Section IV-B. Table VI provides a high-level comparison of surveyed paper using GNNs for spectrum awareness applications. We also provide the dataset or the simulator used for performance evaluation.
TABLE VI: Comparison of surveyed papers using GNN for Spectrum Awareness Applications.
Section Paper Application ML Algorithm Used Dataset/ Simulator Used
Section IV-A Radio Frequency Spectrum Sensing Zhang et al. [105] Cooperative Spectrum Sensing Inductive Graph Neural Network Kriging SPLAT! Tool [215], Terrain database of US Geological Survey, Simulation data [216], Crawdad WiMax [217]
Janu et al. [130] Cooperative Spectrum Sensing Graph Convolution Network Python-based simulator
He et al. [218] Distributed cooperative spectrum sensing GNN + RL Own simulator
Section IV-B Radio frequency signal classification Xuan et al. [123] Modulation classification Adaptive Visibility Graph Neural Network RML2016.10a [219], RML2016.10b [220]
Qiu et al. [221] Radio Signal Classification RNN + CNN + GNN RML2016.10a [219]; own dataset baaed on NVIDIA Tesla T4 simulator (SigData18 and SigData36)
Bertalanič et al. [222] Wireless link layer anomaly classification Graph Isomorphism Network Rutgers WiFi dataset [223]
Pang et al. [224] Network traffic classification Chained Graph Neural Network Application traffic dataset, Malicious traffic dataset, Encrypted traffic dataset [225]
IV-A Radio Frequency Spectrum Sensing
Figure 10: Applications of GNN for RF spectrum sensing tasks.
@@ IV-A Radio Frequency Spectrum Sensing
IV-B Radio Frequency Signal Classification
While detecting the presence of active and inactive spectrum is an important problem, it is equally important to identify the type of signal being transmitted. This is especially important in applications that are critical and vulnerable to external interference. Furthermore, it is important to identify if the signals are friendly or adversary. This classification can usually be done based on the known signals that the incumbent system expects while any other unknown signals can be flagged as an adversary. ML is known for its classification and prediction capabilities across domains. Traditionally, CNNs [226, 227, 228, 229, 230] and RNNs have dominated RF signal classification tasks such as modulation classification, signal type classification, RF Fingerprinting among others [230, 98, 99]. More recently, GNNs are also being leveraged for these tasks including but not limited to [123, 221]. Figure 11 shows common applications of GNN for RF signal classification and prediction tasks. Furthermore, GNN has also impacted other classification tasks such as anomaly classification [222] and traffic classification [224].
AvgNet [123] is an Adaptive Visibility Graph algorithm that maps time series data into graphs and can be used to classify various signal modulations. The main advantage of this algorithm is that it can automatically adapt into best-matching graphs, reducing the time consumed during the mapping process and aggregating both the local and global information of the signal sample. The proposed model is verified using simulation and the results show that AvgNet can effectively classify 11 modulations namely Amplitude Modulation - Double Sideband (AM-DSB), Amplitude Modulation - Single Sideband (AM-SSB), Phase Shift Keying (8PSK), Binary Phase Shift Keying (BPSK), Continuous Phase Frequency Shift Keying (CPFSK), Gaussian Frequency Shift Keying (GFSK), Pulse Amplitude Modulation (PAM4), Quadrature Amplitude Modulation (QAM16, QAM64), Quadrature Phase Shift Keying (QPSK) and Wide-band Frequency Modulation (WBFM).
@@ IV-B Radio Frequency Signal Classification
GNN also finds its application for traffic classification scenarios where it is important to differentiate between traffic generated by various applications while identifying any encrypted or malicious traffic. To enable this, in [224], the authors present CGNN, a Chained GNN to effectively classify various real-world traffic scenarios and categorize them into the different applications as well as label them as “normal”, “encrypted”, or “malicious” traffic. Through extensive simulation, it is shown that compared to CNN and other state-of-the-art techniques, GNN, specifically CGNN outperforms them in terms of three key metrics namely precision, accuracy, and recall.
IV-C Lessons Learned
In this section, we present the key takeaways and lessons learned from the applications of GNN for Spectrum Awareness discussed in Sections IV-A and IV-B.
  1. 1.
    Wireless spectrum environments are inherently complex, featuring multiple users, heterogeneous devices, and dynamic frequency allocations. The interactions between these elements form a structured graph, where nodes represent transmitters, receivers, or frequency bands, and edges capture interference, connectivity, or usage relationships. GNNs excel at learning from graph-structured data, making them well-suited for modeling spectrum environments where traditional machine learning models struggle to account for the intricate dependencies among users and devices. By leveraging message-passing mechanisms, GNNs can infer spectrum availability, predict interference patterns, and optimize spectrum allocation more effectively than conventional approaches.
  2. 2.
    Spectrum occupancy is rarely random; it follows spatiotemporal patterns. Devices in close proximity tend to experience similar spectrum conditions, and usage trends evolve over time due to periodic activities (e.g., office hours, peak traffic times). Traditional models often fail to capture these correlations, leading to suboptimal spectrum predictions. GNNs, however, naturally encode spatial and temporal dependencies by aggregating information from neighboring nodes over multiple graph layers. This enables them to recognize patterns such as congestion hotspots, propagation effects, and historical usage trends, improving spectrum awareness and prediction accuracy.
  3. 3.
    Wireless networks are in constant flux, with devices frequently joining and leaving the network, frequency bands becoming occupied or vacated, and interference levels fluctuating. Static models struggle to adapt to such real-time changes, whereas GNNs, when combined with recurrent or adaptive learning techniques, can dynamically update their representations based on evolving graph structures.
  4. 4.
    Training machine learning models for spectrum awareness often requires labeled datasets, where spectrum occupancy, interference events, and device activities are annotated. However, obtaining labeled data in wireless environments is challenging due to privacy concerns where spectrum usage data may contain sensitive information related to user behavior and communications as well as high data collection costs where Deploying sensors and monitoring devices to label spectrum states is expensive and resource-intensive. GNNs can mitigate data scarcity issues through self-supervised and semi-supervised learning techniques, where they learn from partially labeled graphs or leverage domain knowledge to enhance model generalization. Transfer learning [231, 232] can also be employed, allowing models trained in one spectrum environment to be adapted to another with minimal labeled data.
  5. 5.
    Interference is a major issue in spectrum management, as unauthorized transmissions, overlapping frequency bands, and external noise can degrade network performance. Traditional interference detection methods rely on threshold-based techniques or statistical models, which often fail to capture complex interference patterns arising from diverse sources such as co-channel interference, adjacent-channel leakage, and unintentional electromagnetic interference. GNNs can identify non-trivial interference patterns that traditional models miss. By learning from historical spectrum data and network topology, GNNs can predict interference hotspots, classify interference types, and suggest mitigation strategies such as frequency reallocation or power adjustments.
  6. 6.
    In conclusion, GNNs have shown great potential in improving spectrum awareness, particularly in handling the complex, dynamic, and interconnected nature of wireless networks. As the field evolves, we can expect GNNs to play a critical role in optimizing spectrum use, especially in increasingly crowded and diverse wireless environments.
TABLE VII: Comparison of surveyed papers using GNN for Networking Applications.
Section Paper Application ML Algorithm Used Dataset Used
Section V-A Network and Signal Characteristics Prediction Soto et al. [233] WLAN performance prediction Graph Convolutional Network ITU-T AI Challenge [234], Komondor simulator [235]
Zhou et al. [236] WLAN performance prediction Dynamic Temporal Graph Neural Network Komondor simulator [235]
Rattaro et al. [237] Wireless RSSI prediction Random Dot Product Graphs + GNN Plan Ceibal [238]
Sun et al. [239] Mobile data traffic prediction Graph-based Temporal Convolutional Network Own Wi-Fi dataset
Sun et al. [240] Channel prediction GNN WirelessInSite [241], Blender [242]
Liu et al. [243] Link prediction Graph Capsule Network Six real networks and three citation networks
Section V-B Routing Optimization Yang et al. [134] Joint routing and scheduling optimization GCN + RL NetworkX [244]
Lu et al. [147] Multi-path routing GraphSAGE NS-3 simulator [245]
Rusek et al. [246] Network modeling and optimization GNN OMNET ++ simulator [247]
Almasan et al. [248] Routing optimization GNN + RL OpenAI Gym [249]
Xu et al. [250] Wireless networking routing optimization GNN + RL Own python-based simulator
Swaminathan et al. [251] Routing optimization GNN + RL Own simulator
Yan et al. [252] Multi-path routing GNN Own simulator
Section V-C Congestion Control Kirby et al. [253] Routing congestion prediction Graph Attention Networks Own python-based simulator
LaMar et al. [254] Intent-based routing GNN EXata Network Emulator Software [255]
Ferriol-Galmés et al. [256] Network modeling GNN OMNET-based simulator [247], Routenet-Fermi [257]
Section V-D Mobile Edge Computing Asheralieva et al. [258] Multi-access edge computing GNN + RL OPNET-based simulator [259]
Zeng et al. [260] Distributed edge servers processing GNN SIoT [261], YelpChi [262]
Zhou et al. [263] Network automated design GNN Own python-based simulator
Sun et al. [177] Task offloading GNN + RL Own simulator
Li et al. [264] Edge computing, task offloading GNN + RL Own simulator
Ma et al. [124] Multi-task Offloading Graph Convolutional Network Own simulator
Zhang et al. [265] Service offloading GNN + RL Own python-based simulator
Yang et al. [266] Resource scheduling and Task offloading Graph Convolutional Network Own simulator
Section V-E Digital Twin Networks Wang et al. [267] Network slicing management GNN NSFNET and GEANT2 [268]
Naeem et al. [269] Network Slicing GNN + RL Own simulator
Ferriol-Galmés et al. [270] Network optimization GNN TwinNet [271]
Ferriol-Galmés et al. [272] Network modeling GNN OMNet++ Simulator [247]
Yu et al. [273] Self-healing mechanism in 6G edge networks. GNN OMNet++ based modeling dataset [247]
Zhang et al. [274] Power allocation and user association GNN Own simulator
Section V-F Unmanned Aerial Networks Liu et al. [275] Air Quality Sensing Framework GCN + LSTM Ground and Aerial sensing data [276]
Zhang et al. [277] Cooperative trajectory design Graph Convolutional Neural Network + RL Python-based simulator [278]
An et al. [279] UAV Trajectory Prediction Dynamic GNN Python-based simulator
Mou et al. [280] Self-healing UAV swarm network Graph Convolutional Network Python-based simulator [281]
Wang et al. [282] Joint flying relay location and routing optimization GNN + LSTM Own simulator